OpenStudy (griffindore):

hey i have 2 quick questions if anyone can help

OpenStudy (griffindore):

A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of the following is a possible step used in eliminating the y-term?

OpenStudy (anonymous):

is this type a typed answer?

OpenStudy (griffindore):

(y + z = 6) ⋅ 4 (3y + 4z = 1) ⋅ 4 (y + z = 6) ⋅ −3 (3y + 4z = 1) ⋅ 3

OpenStudy (mathstudent55):

To eliminate a variable, you need to add the equations or multiples of the equations in which the same variable in both equations will add to zero.

OpenStudy (anonymous):

In order to elminiate any term from an equation when you have multiple equations .. U've to multiply one with the first equation that will result in a "zero" when you add them up ... In Ur case, this would be the third choice :D Let me know if U got it :)

OpenStudy (griffindore):

Ok

OpenStudy (griffindore):

oh i get it that makes sence

OpenStudy (mathstudent55):

You want to eliminate y. The second equation has 3y. If you could change the first equation to have -3y, that would do it. Does any choice change the first equation so that the y term becomes -3y?

OpenStudy (anonymous):

sorry I meant to multiply by a constant factor :)

OpenStudy (griffindore):

ya thats what i try but i kept getting stuck

OpenStudy (griffindore):

can u guys help me out with one more

OpenStudy (anonymous):

Feel free :D

OpenStudy (griffindore):

ok

OpenStudy (griffindore):

A set of equations is given below: Equation E: c = 2d + 1 Equation F: c = 3d + 7 Which statement describes a step that can be used to find the solution to the set of equations? Equation F can be written as c = 3(c − 1) + 7. Equation F can be written as c = 2(c − 7) + 1. Equation F can be written as d + 1 = 3d + 7. Equation F can be written as 2d + 1 = 3d + 7.

OpenStudy (anonymous):

Just find d in term of c from Equation E and then plug the value in the other equation or find c in terms of d and do same steps :)

OpenStudy (anonymous):

Let's try to get c in terms of d .. To do so, U've to get c in one hand and d on the other hand in one of the equations let's say equation E .. we already have that done for us .. then plug c in equation F and see if any of the answers is what we have got

OpenStudy (mathstudent55):

Second problem. Look at the given equations. Notice they are both solved for c. Set them equal to each other. If c = 2d + 1 and c = 3d + 7, then 2d + 1 must equal 3d + 7 Which choice shows this?

OpenStudy (anonymous):

If none of them is what we got .. Then we are now sure that the answer contains no d's ! So we try the other way .. Try to get d in terms of c from Equation E and plug it in equation F as well as what we have did in the last step for c in terms of d

OpenStudy (mathstudent55):